A bag contains three red and four black balls. A balls is drawn at random from the black what is the probability that the ball is black?
1
Expert's answer
2020-04-23T16:58:18-0400

P(A​)=Probability of drawing a black ball

N1=4 - number of black balls in the bag

N2=3 - number of red balls in the bag

P(A)=N1/(N1+N2)=4/7=0.57